# Controllable Video Generation: A Survey

### v1

@@ 4.6.1 Text Rendering
Text rendering in the video domain targets synthesizing vivid video with aligned text, which has widespread applications in various forms, including advertisements and movies. Previous works [163, 58, 391, 110, 392, 393] leverage the CLIP [394] text encoder to encode prompts and use 2D VAE [348] to compress the video from RGB to latent space. They fail to generate clear and aligned text. Text-Animator [333] designs a text embedding injection module, a camera control module, and a text refinement module to improve both visual text structures and stability. Recently, Wan [19] has proposed a series of powerful foundation models, including a 1.3B model and a 14B model, which expand the application to various areas such as T2V, text rendering, audio generation, and numerous other downstream tasks.
@@ 4.6.2 Style-Guided Generation
Style-guided video generation aims to transfer an artistic style from a reference image or video to a target video, which is pivotal for creative expression and achieving specific visual aesthetics. The field has progressed from early per-frame stylization with post-hoc temporal smoothing [395, 396] to more efficient and flexible frameworks for arbitrary style transfer [397, 398]. The advent of powerful generative models has marked a significant turning point, with GAN-based approaches enabling specialized applications like video toonification [339] and translation [399]. More recently, diffusion models have become the state-of-the-art, with methods such as Rerender A Video [338] and StyleCrafter [336], which established the viability of diffusion models for high-fidelity video stylization. Subsequent research has expanded this foundation, focusing on greater control and universality. For instance, FRESCO [335] and StyleMaster [334] introduce more refined controls over the stylization process, while UniVST [337] aims to create a more universal framework for handling diverse artistic styles. The latest advancements, including FrameGuidance [89] and OmniVDiff [90], push the boundaries further by introducing novel guidance techniques and omni-controllable capabilities. These state-of-the-art systems commonly utilize mechanisms like cross-attention to inject style information. Despite remarkable progress, ensuring perfect temporal consistency, preserving original content details amidst heavy stylization, and providing intuitive user controls remain active areas of research.
@@ 4.6.3 Point-Guided Generation
Point-guided video generation provides fine-grained control by propagating the influence of sparse user-defined point trajectories to coherently animate specific objects or regions. The central challenge is to naturally extend these sparse constraints over both space and time. Some GAN-based approaches, like Point-to-Point Video Generation [400] and ImaGINator [208], also explore the use of sparse point inputs for controllable video synthesis. Recent methods have introduced effective techniques to address this, often by conditioning powerful pre-trained diffusion models. For example, Drag-a-video [343] pioneers an approach that optimizes latent codes or features to precisely satisfy the point constraints. Another line of work, exemplified by Track4Gen [342], concentrates on robustly tracking user-specified points to ensure that the generated content faithfully adheres to the desired trajectories. Broadening this concept, related frameworks like Diffusion as Shader [340] and GS-DiT [341] show how sparse user input can guide more complex scene properties, hinting at a future of more sophisticated, point-based control. These methods often encode point information as heatmaps or use it to steer attention mechanisms within the network, thereby translating the sparse input into a coherent and dynamic output. Key challenges remain in ensuring physically plausible deformations, maintaining long-term temporal consistency from minimal input, and seamlessly handling complex object interactions and occlusions.
@@ 4.6.4 BEV-Guided Generation
Bird’s-Eye View (BEV) guided video generation is a critical task for creating controllable and realistic driving scenarios for autonomous systems, leveraging a sequence of top-down semantic maps $B=\{b_{1},b_{2},\dots,b_{T}\}$ to dictate scene layout and object dynamics. Earlier works, particularly those based on GANs, explore BEV-conditioned video synthesis methods like DriveGAN [401] for generating realistic driving scenarios. The field rapidly advances, with diffusion models becoming central to translating these abstract BEV representations into photorealistic videos. Foundational works like MagicDrive [123] and its successor MagicDrive-V2 [124] establish high-fidelity BEV-conditioned video synthesis. Subsequent research significantly expands these capabilities. For instance, DriveDreamer-2 [122] focuses on creating interactive simulations, while DiVE [119] and Delphi [120] aim to generate diverse and realistic driving environments. Concurrently, methods like DreamForge [118] work towards building high-fidelity world models, Panacea [127] introduces novel controllable generation techniques, and Seeing Beyond Views [344] tackles the challenge of generating novel perspectives from limited inputs. Despite this progress, significant challenges persist, including ensuring strict geometric consistency between the BEV input and the rendered video, achieving photorealism in object appearance and lighting, accurately simulating complex dynamic interactions, and bridging the semantic gap between the abstract BEV data and the fine-grained details of the visual world.
@@ 4.7 Universal Control
Universal-guided control refers to the ability to flexibly manage and integrate various types of input conditions—such as text, spatial features, and temporal information—to generate videos that meet specific user requirements. This method allows for high customization, enabling users to influence multiple aspects of the generated video simultaneously, such as object behavior, motion, camera angles, and scene layout. By unifying different conditions into a single, coherent framework, universal control facilitates comprehensive video synthesis with enhanced flexibility and precision.
@@
In universal-guided video generation, the model typically incorporates a range of inputs, including text, spatial conditions (such as camera poses, depth maps, and human poses), and temporal signals (including motion sequences and video clips). For example, frameworks like VideoComposer [345] utilize motion vectors or sketches as explicit control signals, guiding the video generation process. Similarly, Any2Caption [346] employs multimodal inputs, such as text, camera poses, and human motions, which are interpreted into structured captions to drive video synthesis. FullDiT [347] integrates these various conditions using a unified attention mechanism, processing text, camera views, depth information, and identity conditions simultaneously, thus enabling seamless video generation.
The methods for universal control in video generation are primarily based on autoregressive and diffusion models. Autoregressive methods, such as those employed in Any2Caption, interpret various conditions (e.g., motion, style, camera) and generate structured captions that can be fed into video generators. These systems usually rely on transformer architectures, such as the one used in FullDiT, which enables multi-task learning with full attention mechanisms, accommodating multiple inputs efficiently.
On the other hand, diffusion models like VideoComposer[345] employ a video latent diffusion model (VLDM) where the video generation process is conditioned by various spatial and temporal conditions. VideoComposer specifically uses a spatio-temporal condition encoder (STC-encoder) that integrates the input conditions, ensuring that motion, camera, and textual elements are synthesized cohesively. This approach is highly flexible, allowing for the dynamic adaptation of diverse input types without requiring fine-tuning.
@@
Despite the progress made in universal-guided control, several challenges persist. One of the primary difficulties lies in balancing the integration of diverse conditions. Ensuring that each condition (e.g., motion, text, camera) is harmoniously combined without one overshadowing the others remains a complex task. In addition, maintaining temporal consistency across frames presents another challenge, particularly when dealing with dynamic and varied motion patterns. Furthermore, evaluation continues to be an unresolved issue—while benchmarks like FullBench for FullDiT have been introduced, the development of comprehensive and standardized evaluation criteria for universal control across all modalities is still underway.
@@
Figure 5: Applications of Controllable Video Generation. The sample images are from DiffuEraser [402], VideoAnydoor [403], 4D-fy [404], Vista [405], spmem [406], RoboMaster [407].
@@ 5 Application
With the rapid advancement of video generative models, video generation techniques are increasingly being applied to real-world scenarios. As illustrated in Fig. 5, T2V models have demonstrated significant potential in areas such as video inpainting, object composition, 4D generation, autonomous driving, world model, and embodied intelligence.
5.1 Video Completion and Inpainting
@@
Video completion and inpainting aim to fill in missing areas in videos or remove unnecessary objects while maintaining spatial and temporal consistency. Traditional methods often struggle with large occlusions or complex motion. In recent years, video generative models have shown strong generative capabilities in this field. Based on the stable diffusion model, DiffuEraser [402] is designed to fill occluded areas, combining prior information for initialization and weak conditional constraints to help reduce noise artifacts and suppress hallucination. AVID [408] introduces motion modules and adjustable structural guidance, which support different types of repair under various levels of structural fidelity, capable of handling videos of any length, ensuring temporal consistency within the editing area.
5.2 Video Composition
@@
Video composition, particularly object insertion, aims to seamlessly embed target objects into existing videos while ensuring both temporal and spatial consistency. This task requires the inserted object to be well integrated with the background scene in terms of motion, lighting, and style while preserving the details of the appearance of the object. Recent advances in text-to-video diffusion models offer powerful control capabilities for this purpose. For example, MVOC [409] reverses the corresponding noise features by DDIM inversion on each video object, then synthesizes and edits them to generate the first frame of the synthetic video. VideoAnydoor [403] is an end-to-end video object insertion framework that achieves high-fidelity detail retention and precise motion control through the ID extractor, Pixel Warper, and optimized training strategies.
5.3 Video-to-4D Generation
@@
4D generation is more complex in both training and modeling compared to static 3D generation. Video generative models can serve as motion priors, effectively capturing scene dynamics, and emerging as a core technology for constructing 4D representations. 4D-fy [404] proposes a mixed score distillation method that combines multiple pre-trained diffusion models to generate 4D scenes with realistic appearance, structure, and motion. 4Real [410] utilizes a video diffusion model to generate reference videos and “freeze-time” videos, and learns a standard 3D representation based on deformable 3D Gaussian splatting and time-varying deformations from them.
5.4 Autonomous Vehicle and World Generative Model
@@
World models play a vital role in autonomous driving by simulating environments and guiding decision-making. Video generation models, with their ability to produce realistic visual data, are now increasingly used to improve perception and planning in autonomous systems. GAIA-2 [411] is a multi-view generative world model capable of producing high-resolution, temporally and spatially consistent multi-camera videos from structured inputs such as vehicle dynamics, environmental conditions, and road semantics. DriVerse [412] generates high-fidelity driving simulation videos by using a single image and future trajectories, introduces multi-modal trajectory prompts to encode trajectories as text and spatial motion priors, and uses latent motion alignment to enhance the temporal consistency of dynamic objects. Drive-WM [413] promotes spatial-temporal joint modeling through view factorization, enabling the generation of high-fidelity multi-view videos in autonomous driving scenarios. Based on its powerful generative capability, this method demonstrates the potential of applying world models to safe driving planning.
5.5 Embodied Artificial Intelligence
@@
Video generative models can simulate future scenarios, enabling embodied intelligent agents to understand, predict, and plan interactive behaviors by generating visual information that aligns with semantic and physical constraints. This enhances the agents’ capabilities within the perception–decision–action loop. Gen2Act [414] introduces a zero-shot framework that generates human operation videos using a video generation model trained on large-scale internet data. These videos are then translated into executable robot actions via a policy model, addressing the challenge of generalizing to unseen objects and novel tasks. AVDC [415] leverages a video diffusion model to synthesize hallucinated videos of robots performing actions. It extracts dense correspondences via optical flow within the synthesized video and combines them with initial depth estimates to regress rigid transformations of the target scene or object, thereby inferring robot actions. RoboMaster [407] generates robot manipulation videos through collaborative trajectory control. It addresses the feature entanglement problem in multi-object interaction by decomposing the process into three stages and integrating object embeddings with appearance and shape perception. This&That [416] enhances robot planning and execution by generating task videos conditioned on both language and gestures. Compared to language-only methods, this multimodal approach enables more accurate interpretation of user intentions.
@@ 6 Discussion and Future Work
The field of controllable video generation has made remarkable progress, as evidenced by the representative works summarized in Tab. II, yet several critical challenges remain unaddressed. These challenges hinder the development of robust, scalable, and universally controllable video generation systems. To help future research, in this section, we discuss three potential research directions.
6.1 Unified Control Mechanisms
Achieving unified control in video generation poses a significant challenge due to the need to balance multiple constraints, such as temporal consistency, spatial fidelity, motion dynamics, and stylistic coherence. Current methods often address these aspects in isolation, leading to inevitable trade-offs. For example, structure control methods like Follow-Your-Pose [2] and DriveDreamer [122] excel at synthesizing videos based on specific input structures, such as poses or bounding boxes. Similarly, temporal control approaches like MotionBooth [152] and Direct-A-Video [220] focus on maintaining temporal consistency. Meanwhile, image control methods such as VideoCrafter1 [3] and Lumiere [176] achieve impressive results in style transfer and image-guided synthesis.
To address these limitations, future research should focus on developing hierarchical control frameworks that enable the compositional integration of constraints. These frameworks could dynamically prioritize user-defined controls based on task complexity. Another promising direction involves designing adaptive mechanisms that respond to user preferences in real-time, enabling iterative refinement of constraints during the generation process. Adaptive systems could prioritize temporal consistency in scenarios requiring smooth action sequences but shift focus to stylistic attributes when generating animated or artistic video content. Early works like VideoComposer [345] represent a step toward such adaptability by supporting universal conditions, but further refinement is required to handle diverse and complex user inputs effectively.
Additionally, compositional diffusion approaches—where generation tasks are broken into modular stages—could significantly enhance the flexibility of video generation systems. For instance, models like FullDiT [347] demonstrate the potential of integrating multiple conditions (e.g., text, image, and motion) into a unified framework. However, these approaches often face challenges in balancing trade-offs between conditions and ensuring computational efficiency. Future efforts should aim to design scalable frameworks that seamlessly integrate multimodal conditions while maintaining high-quality outputs.
6.2 Unified Video Reasoning + Generation
Large Language Models (LLMs) have demonstrated remarkable potential in understanding user intent and enabling multi-modal reasoning, making them essential for advancing video generation and controllability. By integrating LLMs with video generative models, users can interact with these systems using natural language prompts to specify tasks or refine outputs iteratively. For instance, frameworks like SkyReels-A2 [128] and Phantom [142] show how textual inputs, combined with other modalities such as object details or scene descriptions, can produce semantically rich videos. However, challenges remain in aligning LLMs with video generation systems to handle complex, multi-turn interactions.
One promising direction for future research is the development of multi-modal alignment frameworks that bridge the gap between LLMs and video generative models. For example, works like VideoCrafter1[3] have shown early progress in combining text and image conditions for video generation. Expanding these approaches by leveraging LLMs for cross-modal understanding could enable users to combine diverse inputs such as sketches, reference images, or audio cues alongside text to guide the generation process.
In this context, Multi-Modal LLM (MLLM) agents are emerging as transformative tools for video generation and reasoning. These agents extend the capabilities of traditional LLMs by processing and responding to inputs across diverse formats, such as text, images, audio, and video. Acting as a unified interface, MLLM agents simplify complex workflows by reasoning across modalities. For example, an MLLM agent could process a text description like, “Create a video of a futuristic cityscape with flying cars,” alongside a reference sketch of the city layout and an audio clip of background sound effects. The agent would integrate these inputs, reason across modalities, and generate a coherent video aligned with the user’s vision. They also excel in adaptive refinement: users could iteratively interact with the agent to adjust outputs, offering feedback like, “Make the cars more sleek and futuristic,” or, “Adjust the lighting to a golden sunset.” These capabilities make MLLM agents invaluable for tasks such as storyboarding, cinematic planning, and educational video creation.
Addressing challenges like bias mitigation and computational efficiency is vital to make these systems practical and responsible. Bias detection and correction techniques are necessary to avoid unintended outputs, while lightweight fine-tuning and efficient cross-modal adapters can reduce computational overhead. By aligning LLMs with video generative models and harnessing MLLM agents, future systems could unlock new possibilities for high-quality, semantically adaptive, and user-driven video generation tailored to diverse needs.

### v2

@@ 5.6.1 Text Rendering
Text rendering in the video domain targets synthesizing vivid video with aligned text, which has widespread applications in various forms, including advertisements and movies. Previous works [jiang2023videobooth, shi2024fonts, shi2025wordcon, zeng2023makepixelsdancehighdynamic] leverage the CLIP [CLIP] text encoder to encode prompts and use 2D VAE [jiang2025vace] to compress the video from RGB to latent space. They fail to generate clear and aligned text. Text-Animator [liu2024textanimatorcontrollablevisualtext] designs a text embedding injection module, a camera control module, and a text refinement module to improve both visual text structures and stability. Recently, Wan [wan2025] has proposed a series of powerful foundation models, including a 1.3B model and a 14B model, which expand the application to various areas such as T2V, text rendering, audio generation, and numerous other downstream tasks.
@@ 5.6.2 Style-Guided Generation
Style-guided video generation transfers an artistic style from a reference image or video to a target video, which is central for creative expression and for achieving specific visual aesthetics. The field has progressed from early per-frame stylization with post-hoc temporal smoothing [chen2017coherent, huang2017real] to efficient and flexible frameworks for arbitrary style transfer [song2025omniconsistency, guo2025any2anytryon, song2024diffsim, song2024processpainter, zhang2025stable, chen2025transanimate, zhang2024stable, deng2021mccnet, gao2020fast]. With the development of generative models, GAN-based approaches support specialized applications such as video toonification [yang2022vtoonify] and video translation [yang2023styleganex], while more recent diffusion models, including Rerender A Video [yang2023rerender] and StyleCrafter [liu2023stylecrafter], now form the leading approach for high-fidelity video stylization. Building on this, FRESCO [yang2024fresco] and StyleMaster [ye2024stylemaster] introduce more fine-grained stylization controls, while UniVST [song2024univst] aims at a general framework for diverse artistic styles. Remaining challenges include temporal consistency, preservation of content details under strong stylization, and intuitive user controls.
@@ 5.6.3 Point-Guided Generation
Point-guided video generation provides fine-grained control by using sparse user-defined point trajectories to animate specific objects or regions, and the main challenge is to extend these constraints in a stable way over space and time. GAN-based methods such as Point-to-Point Video Generation [wang2019point] and ImaGINator [wang2020imaginator] first explore sparse point inputs for controllable video synthesis, while later work often conditions pre-trained diffusion models on such inputs. Track4Gen [jeong2024track4gen] focuses on robustly tracking user-specified points so that the generated content follows the desired trajectories, and related frameworks such as Diffusion as Shader [gu2025diffusionasshader] and GS-DiT [bian2025gs] show that sparse user input can steer more complex scene properties. These methods typically encode points as heatmaps or use them to guide attention mechanisms, translating sparse input into coherent and dynamic outputs. Key challenges include physically plausible deformations, long-term temporal consistency from minimal input, and handling complex object interactions and occlusions.
@@ 5.6.4 BEV-Guided Generation
Bird’s-Eye View (BEV) guided video generation targets controllable and realistic driving scenarios for autonomous systems, where a sequence of top-down semantic maps $B=\{b_{1},b_{2},\dots,b_{T}\}$ specifies scene layout and object dynamics. Early GAN-based methods such as DriveGAN [kim2021drivegan] study BEV-conditioned video synthesis for realistic driving sequences, and recent work establishes diffusion models as a central paradigm for translating abstract BEV representations into photorealistic videos, as illustrated by MagicDrive [gao2023magicdrive], MagicDrive-V2 [gao2024magicdrivedit], DriveDreamer-2 [zhao2025drivedreamer], DiVE [jiang2024dive], Delphi [ma2024unleashing], DreamForge [mei2024dreamforge], Panacea [wen2024panacea], and Seeing Beyond Views [lu2024seeing]. Despite this progress, key challenges remain in geometric consistency between BEV inputs and rendered videos, realistic appearance and lighting, accurate modeling of complex dynamic interactions, and the semantic gap between abstract BEV representations and fine-grained visual details.
5.6.5 Summary
Other control modalities, including text rendering, style transfer, point tracking, and BEV conditioning, provide fine-grained and semantically aligned control over video generation beyond identity, structure, and motion. In our framework, these modalities connect foundation models with high-level creative tools through conditional encoders and control modules such as adapters and cross-attention mechanisms, which inject sparse or abstract guidance into the generative process. The design trend is shifting from mainly UNet-based architectures (for example, ControlNet-based stylization) toward DiT-based backbones, as in Wan and advanced BEV models, to better capture global semantic context and complex spatio-temporal dependencies for tasks such as coherent text spelling and autonomous driving simulation. From a user perspective, these controls support personalization through style and text and real-time preview and adjustment through point dragging, while turning these mechanisms into intuitive, transparent tools with effective automatic suggestions remains an important direction for future research.
@@ 5.7 Universal Control
Universal Control refers to the ability to flexibly manage and integrate various types of input conditions—such as text, spatial features, and temporal information—to generate videos that meet specific user requirements. This method allows for high customization, enabling users to influence multiple aspects of the generated video simultaneously, such as object behavior, motion, camera angles, and scene layout. By unifying different conditions into a single, coherent framework, universal control facilitates comprehensive video synthesis with enhanced flexibility and precision.
@@
In universal-guided video generation, the model typically incorporates a range of inputs, including text, spatial conditions (such as camera poses, depth maps, and human poses), and temporal signals (including motion sequences and video clips). For example, frameworks like VideoComposer [wang2023videocomposer] utilize motion vectors or sketches as explicit control signals, guiding the video generation process. Similarly, Any2Caption [wu2025any2caption] employs multimodal inputs, such as text, camera poses, and human motions, which are interpreted into structured captions to drive video synthesis. FullDiT [ju2025fulldit] integrates these various conditions using a unified attention mechanism, processing text, camera views, depth information, and identity conditions simultaneously, thus enabling seamless video generation.
Several recent frameworks highlight different ways to achieve this. For instance, VideoComposer [wang2023videocomposer] relies on a dedicated spatio-temporal condition encoder to blend motion cues, depth information, and text prompts inside a latent diffusion pipeline. Any2Caption [wu2025any2caption], on the other hand, takes a more indirect route: it first translates diverse inputs into detailed, structured textual captions using a multimodal LLM, then feeds those captions to a standard video generator. FullDiT [ju2025fulldit] pushes direct integration further by employing full spatiotemporal attention within a Diffusion Transformer architecture, letting it handle text, camera poses, depth, and even identity constraints all at once. Notably, in the autoregressive camp, InfinityStar [liu2025infinitystar] stands out—it simply tokenizes everything (visual content and controls alike) into a unified discrete sequence, enabling native multi-condition control for tasks ranging from text-to-video to image animation and long-sequence extension, all without extra adapters.
@@
Despite the progress made in universal-guided control, several challenges persist. One of the primary difficulties lies in balancing the integration of various conditions. Ensuring that each condition (e.g., motion, text, camera) is harmoniously combined without one overshadowing the others remains a complex task. In addition, maintaining temporal consistency across frames presents another challenge, particularly when dealing with dynamic and varied motion patterns. Furthermore, evaluation continues to be an unresolved issue—while benchmarks like FullBench for FullDiT have been introduced, the development of comprehensive and standardized evaluation criteria for universal control across all modalities is still underway.
In summary, universal control uses advanced conditioning—from encoders and attention fusion to token unification—to combine signals like text, poses, depth, and motion. This lets users customize content, style, motion, and layout all at once, far surpassing what single-condition models can do.
The architecture has shifted from bulky UNet-with-adapters designs to more scalable backbones, driven by the need for cleaner multi-condition handling and better scaling. Diffusion Transformers (e.g., FullDiT [ju2025fulldit]) and autoregressive models (e.g., InfinityStar [liu2025infinitystar]) now lead, offering natural fusion of controls, support for long/high-res sequences, and either high fidelity or fast in-context flexibility. Caption approaches like Any2Caption [wu2025any2caption] and early composable encoders from VideoComposer [wang2023videocomposer] remain useful, but trends favor integrated, interactive systems. Challenges in signal conflicts, long-term coherence, and interpretability persist, suggesting future progress will focus on hybrid designs and clearer user insight into control interactions.
@@
Figure 7: Applications of Controllable Video Generation. The sample images are from DiffuEraser [li2025diffueraserdiffusionmodelvideo], VideoAnydoor [tu2025videoanydoorhighfidelityvideoobject], 4D-fy [4d-fy], Vista [gao2024Vista], spmem [wu2025spmem], RoboMaster [fu2025learningvideogenerationrobotic].
@@ 6 Application
With the rapid advancement of video generative models, video generation techniques are increasingly being applied to real-world scenarios. As illustrated in Fig. 7, T2V models have demonstrated significant potential in various areas.
@@
Video Completion and Inpainting. Video completion and inpainting aim to fill in missing areas in videos or remove unnecessary objects while maintaining spatial and temporal consistency. The techniques can be widely used in film post-production, historical video restoration, and everyday mobile video editing apps. Traditional methods often struggle with large occlusions or complex motion. In recent years, video generative models have shown strong generative capabilities in this field. DiffuEraser [li2025diffueraserdiffusionmodelvideo] is designed to fill occluded areas, combining prior information for initialization and weak conditional constraints to help reduce noise artifacts and suppress hallucination. AVID [zhang2024avidanylengthvideoinpainting] introduces motion modules and adjustable structural guidance, which support different types of repair under various levels of structural fidelity, capable of handling videos of any length, ensuring temporal consistency within the editing area.
@@
Video Composition. Video composition aims to seamlessly embed target objects into existing videos while ensuring both temporal and spatial consistency. It is essential in advertising production, AR/VR environments, and creator tools such as short-video platforms. It allows users to insert virtual objects into real scenes, enabling applications like virtual product placement, AR furniture try-on, or creative storytelling with animated characters. It significantly reduces the effort needed to produce high-quality mixed-reality content. This task requires the inserted object to be well integrated with the background while preserving the details of the appearance of object. Recent advances in text-to-video diffusion models offer powerful control capabilities for this purpose. For example, MVOC [wang2024mvoctrainingfreemultiplevideo] reverses the corresponding noise features by DDIM inversion on each video object, then synthesizes and edits them to generate the first frame of the synthetic video. VideoAnydoor [tu2025videoanydoorhighfidelityvideoobject] is an end-to-end video object insertion framework that achieves high-fidelity detail retention and precise motion control through the ID extractor, Pixel Warper, and optimized training strategies.
@@
Video-to-4D Generation. 4D generation is inherently more challenging than static 3D reconstruction, as it requires modeling both spatial geometry and temporal dynamics. Video generative models naturally provide strong motion priors by capturing scene dynamics over time, making them an emerging core technology for constructing high-quality 4D representations. It enables game studios, film creators, and cultural heritage organizations to build 4D digital assets from simple handheld videos, avoiding expensive multi-camera capture systems. It can also support robotic perception by providing realistic, dynamic 3D environments for simulation and training. 4D-fy [4d-fy] proposes a mixed score distillation method that combines multiple pre-trained diffusion models to generate 4D scenes with realistic appearance, structure, and motion. 4Real [4real] utilizes a video diffusion model to generate reference videos and “freeze-time” videos, and learns a standard 3D representation based on deformable 3D Gaussian splatting and time-varying deformations.
@@
Autonomous Vehicle and Generative World Model. World model is important in autonomous driving by simulating environments and guiding decision-making. Video generation models, with their ability to produce realistic visual data, are now increasingly used to improve perception and planning in autonomous systems. GAIA-2 [GAIA-2] is a multi-view generative world model capable of producing high-resolution, temporally and spatially consistent multi-camera videos from structured inputs such as vehicle dynamics, environmental conditions, and road semantics. DriVerse [DriVerse] generates high-fidelity driving simulation videos by using a single image and future trajectories, introduces multi-modal trajectory prompts to encode trajectories as text and spatial motion priors, and uses latent motion alignment to enhance the temporal consistency of dynamic objects. Drive-WM [wang2023Drivingintothefuture] promotes spatial-temporal joint modeling through view factorization, enabling the generation of high-fidelity multi-view videos in autonomous driving scenarios. The powerful generative capability demonstrates the potential of applying world models to safe driving planning.
@@
Embodied Artificial Intelligence. Video generative models can simulate future scenarios, enabling embodied agents to understand, predict, and plan interactive behaviors by generating visual information that aligns with semantic and physical constraints. This enhances the agents’ capabilities within the perception–decision–action loop. Gen2Act [Gen2Act] introduces a framework that uses video generation model to generate human operation videos. These videos are translated into executable robot actions via a policy model, addressing the challenge of generalizing to unseen objects and novel tasks. AVDC [AVDC] leverages video diffusion model to synthesize hallucinated videos of robots performing actions. It extracts dense correspondences via optical flow within the synthesized video and combines them with initial depth estimates, thereby inferring robot actions. RoboMaster [fu2025learningvideogenerationrobotic] generates robot manipulation videos through collaborative trajectory control. It addresses the feature entanglement problem in multi-object interaction by decomposing the process into three stages and integrating object embeddings with appearance and shape perception. This&That [wang2024language] enhances robot planning and execution by generating task videos conditioned on both language and gestures. Compared to language-only methods, this multimodal approach enables more accurate interpretation of user intentions.
@@ 7 Discussion and Future Work
Controllable video generation has advanced rapidly (Tab. 2), yet building robust and scalable systems that support diverse controls remains challenging. Current methods often optimize a subset of constraints (e.g., structure, temporal consistency, style), leading to unavoidable trade-offs when multiple conditions must be satisfied simultaneously. We highlight three promising directions for future research.
7.1 Unified Multi-Condition Control and Constraint Orchestration
A key bottleneck is compositional controllability: real applications usually require jointly satisfying temporal coherence, spatial/structural fidelity, motion dynamics, and stylistic consistency. Existing works typically emphasize one aspect at a time—e.g., structure-driven generation such as Follow-Your-Pose [ma2024followposeposeguidedtexttovideo] and DriveDreamer [zhao2025drivedreamer], temporal control such as MotionBooth [wu2024motionbooth] and Direct-A-Video [yang2024directavideo], and image/style-guided generation such as VideoCrafter1 [chen2023videocrafter1opendiffusionmodels] and Lumiere [10.1145/3680528.3687614]. As a result, enforcing one constraint often degrades others.
Future systems should move toward constraint orchestration: (i) hierarchical or modular control frameworks that compose heterogeneous conditions and explicitly manage conflicts; (ii) adaptive control policies that can re-weight constraints based on user intent and content (e.g., prioritizing temporal smoothness for action sequences while emphasizing style for artistic videos). VideoComposer [wang2023videocomposer] is an early step toward universal conditioning, but stronger mechanisms are needed to reliably handle complex, mixed user inputs. Compositional diffusion and unified architectures (e.g., FullDiT [ju2025fulldit]) further suggest the feasibility of multi-condition integration, yet open problems remain in balancing competing constraints and maintaining computational efficiency.
7.2 Unified Video Reasoning and Generation
Another direction is to couple reasoning with generation [chen2025videochat, chen2025top, chen2025g]. LLMs can serve as a semantic control layer that translates high-level intent into concrete, low-level conditions and supports iterative refinement through multi-turn interaction. Recent systems such as Phantom [liu2025phantom] demonstrate progress toward semantically rich, prompt-driven generation, but robust alignment between reasoning modules and video generators is still limited, especially under complex instructions and long-horizon consistency requirements.
A promising route is multi-modal alignment that allows users to combine text with other signals while keeping the interaction natural and editable. In this context, Multi-Modal LLM (MLLM) agents can act as a unified interface for planning, grounding, and refinement: they can parse intent, select or synthesize appropriate controls, and iteratively update constraints based on user feedback. Key challenges include (i) reliable grounding and verification to prevent instruction drift and hallucinated content; (ii) bias and safety mitigation in open-ended generation; and (iii) efficiency, e.g., lightweight adapters or cross-modal modules that reduce compute overhead while preserving controllability.